# Spurious reconstruction from brain activity

### v3

@@ Discussion
Additionally, the independence of test stimuli from the model training processes requires careful examination, particularly when using pre-trained deep neural network (DNN) models and foundation models like CLIP or diffusion models. These models are typically trained on vast amounts of data available on the internet [22, 24, 50], which likely includes public datasets such as MS-COCO [94] that are used in the NSD. This overlap raises potential concerns about the true independence of test stimuli, as we cannot rule out the possibility that these pre-trained models have acquired representations specifically tailored to the stimuli used for the model training. To address these issues and ensure a more rigorous evaluation of model generalization capabilities, researchers may consider using test stimuli that are not publicly available on the internet. This could include self-created stimuli or carefully curated datasets that have not been used in the training of widely used AI models [29, 32]. Such an approach would provide a more stringent test of a model’s ability to generalize to truly novel inputs.
@@
One of the remaining challenges in visual image reconstruction is the development of metrics for evaluating the quality and accuracy of the reconstructed images. The first and most critical step in assessing reconstruction results is to confirm a qualitative similarity between the reconstructed images and the perceived images through visual inspection across a diverse range of test sets. Following this, quantitative metrics should be employed for a more objective, high-throughput evaluation. However, as our analysis has suggested, it can be misleading to evaluate reconstruction by heavily relying on identification performance based on the relative similarity among alternatives [58]. Even in cases where the reconstructed images only capture superficial information, such as categories or overall brightness, identification metrics can still be high. While identification performance can provide a useful benchmark, it should not be the sole metric for evaluating reconstruction quality. It is crucial to develop more appropriate similarity metrics that can accurately measure the perceptual similarity between the reconstructed and original images. One promising approach is to leverage image quality assessment (IQA) techniques from the computer vision field [55, 95]. These techniques are designed to quantify the perceptual quality of images and can be adapted to the specific requirements of visual image reconstruction.
@@
Our findings have significant implications for research integrity and responsible dissemination of scientific results. Visual image reconstruction methods have gained attention not only from neuroscientists but also from the general public and policymakers, sparking discussions about their potential applications and risks [8]. These stakeholders often contemplate the possibilities of seamless information communication through the brain, such as in brain-machine interfaces (BMIs), or the dangers of unauthorized access to private information from brain activity. This interest may stem from the perception that brain activity data can be obtained easily and reliably in real-time. However, current technology and analysis methods fall short of these expectations. Beyond the limitations discussed in this study, there are additional challenges in the field. Most reconstruction methods analyze previously acquired brain data offline. The brain data used for reconstructing images are often averaged over multiple presentations of the test image, with only a few studies demonstrating single-trial reconstruction results [28, 32]. Further, it has been argued that subject cooperation is essential for reliably training and testing decoding models [71]. Given these realities, public expectations often exceed current capabilities, and meeting these high demands in the short term is challenging. By clearly articulating these constraints, we can help manage expectations, prevent disappointment, and guide governments and companies away from misguided decisions. It is crucial to resist making overly optimistic claims about the ability to reconstruct arbitrary images. In light of these challenges, while the field of visual image reconstruction from brain activity holds great promise, it is our responsibility as researchers to ensure that its current capabilities and limitations are accurately communicated to all stakeholders.
Materials and methods
@@
We utilized two datasets: the Natural Scene Dataset (NSD) [12] and the Deeprecon dataset [29]. Both datasets comprise visual stimuli and corresponding fMRI activity collected when subjects perceived the stimuli. In the NSD dataset, eight subjects were presented with MSCOCO images [94], yielding $30,000$ brain activity samples per subject, which is three times the amount provided by the Deeprecon dataset. The Deeprecon dataset includes fMRI activity data from subjects presented with both ImageNet images [9] and artificial images. It contains roughly $8,000$ brain samples per subject. Since this dataset is designed to evaluate reconstruction performance, the test stimuli were carefully selected. The test natural images were selected from ImageNet, which were in categories different from those used in the training. The artificial images were only used as test data to check the generalizability performance of the proposed reconstruction methods. In both datasets, we adopted the train-test split used in previous studies and utilized data from the first subject (S1 in the NSD and Subject 1 in the Deeprecon). Text-guided reconstruction methods require text annotations of images. For the NSD, text annotations accompanying the MSCOCO database were used. For the Deeprecon dataset, we collected text annotations for each experimental stimulus via crowd workers on Amazon Mechanical Turk, yielding five annotations per image. The text annotations of training stimuli are publicly available in the GitHub repository (https://github.com/KamitaniLab/GOD_stimuli_annotations).
@@
We utilized three image reconstruction methods: StableDiffusionReconstruction [44], Brain-Diffuser [45], and iCNN [29]. Each method employs two common steps: first, translating brain activity patterns into latent features of the stimuli, and second, generating images from these latent features using an image generator (see also Fig 1). In the StableDiffusionReconstruction method, the latent features are the VAE [51] features calculated from stimulus images and the CLIP text features [22] from the image annotations. The generator is StableDiffusion [50]. They first generate low-resolution images from the translated VAE features, and those images are further fed into the StableDiffusion model with translated text features to generate images. The generated images are regarded as reconstructed images from brain activity. In the Brain-Diffuser method, the latent features are the VDVAE [52], CLIP vision features from stimulus images and CLIP text features from the text annotations of the image. The generator is Versatile diffusion [49]. Similar to StableDiffusionReconstruction, low-resolution images are first generated from the translated VDVAE features, and these images are further used for the input of the versatile diffusion model with the translated vision and text features. The generated images are regarded as reconstructed images from brain activity. In the iCNN method, the latent features are the intermediate output of the VGG19 layer [56]. As a generator, they used the pre-trained image generator [69], and they solved the optimization problem to minimize the discrepancy between the VGG19 features calculated from the generated images and the translated VGG19 features. Well-optimized images are regarded as reconstructed images.
@@
To investigate dataset diversity, we employed uniform manifold approximation and projection (UMAP), a non-linear dimensionality reduction technique [54], to learn a projection from a latent features space to a lower dimension (UMAP embedding space). We used both the training and test CLIP text features to learn the UMAP projection. These features were combined and standardized beforehand. The UMAP hyperparameters followed the official UMAP guide for clustering usage (https://umap-learn.readthedocs.io/en/latest/clustering.html#umap-enhanced-clustering) with cosine distance as a distance metric (the visualization result with the default UMAP parameters is shown in S1 Fig.). The learned UMAP was also used to project the features predicted from brain activity (Fig 6B). After standardizing the predicted features using the same mean and standard deviation parameters used in UMAP projection learning, we projected the predicted features onto the UMAP embedding space.
@@
We conducted a simulation analysis to illustrate output dimension collapse (Fig 8) and to examine the generalization performance beyond the training data (Fig 9). These analyses involve a teacher-student learning task. The teacher model generated the pairs of target features and brain activity as training samples, and the student model learned a mapping from the training samples. To imitate the feature translation situation from brain activity, observation noise was added to the brain data. To imitate the feature translation situation from brain activity, observation noise was added to the brain data.
@@
Using $N$ training samples, $X_{\mathrm{tr}}=[\mathbf{x}^{(1)},\ldots,\mathbf{x}^{(N)}]^{\top}$ and $Y_{\mathrm{tr}}=[\mathbf{y}^{(1)},\ldots,\mathbf{y}^{(N)}]^{\top}$, we trained the student model. by the ridge regression algorithm. The trained weight of ridge regression model $W$ can be calculated analytically:
@@
After training the student weight $W$, we illustrated the phenomenon of output dimension collapse as in Fig 8 by predicting randomly generated data. First, we generated random test latent features from the Gaussian distribution with a mean of $\mathbf{0}$ and a variance equal to the training set variance scaled by a single scalar value ($9.0$). The corresponding test brain samples were obtained by translating the latent features by the teacher’s weight $\bar{A}$ and adding observation noise $\boldsymbol{\xi}$. The predicted latent features were derived by projecting the test brain samples by the learned student weight $W$. We set $D=512$, $\sigma_{\mathrm{intra}}^{2}=10/512$, $\sigma_{\mathrm{inter}}^{2}=100/512$, and $\sigma_{\mathrm{noise}}^{2}=10$.
We evaluated the zero-shot performance of the student model as in Fig 9. We prepared two types of test samples: in-distribution test samples and out-of-distribution test samples. In-distribution test samples are generated from one of the (Gaussian) clusters used in training data:
@@
We mainly focused on the typical cluster identification performance for dimension $D$, the number of clusters in the training set $C$, and the ratio of variance about the cluster structure $\sigma_{\mathrm{intra}}^{2}/\sigma_{\mathrm{inter}}^{2}$. We used a sufficiently large number of training samples $N=500,000$. This size was constant when the above parameters were varied. Other parameters were also fixed in this simulation as follows: $\sigma_{\mathrm{intra}}^{2}+\sigma_{\mathrm{inter}}^{2}=110/D$, ${\bar{A}_{ij}}\sim\mathcal{N}(0,D^{-1/2})$, $\sigma_{\mathrm{noise}}^{2}=0.25$, $\lambda=1.0$, $n=100$ and $t=32$. We parameterized the scale of the variances ($\sigma_{\mathrm{intra}}^{2}$ and $\sigma_{\mathrm{inter}}^{2}$) and the teacher weights $\bar{A}_{ij}$ using the dimension $D$ so that the order of scale of the output samples were invariant from the dimension. When the number of clusters in the training set $C$ was less than $t=32$, we set $t=C$ instead.
@@
$\mathrm{Acc}$ $\underset{\phantom{n\to\infty}}{=}\frac{1}{n(n-1)}\left(\sum_{(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in Z}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)+\frac{1}{n(n-1)}\left(\sum_{(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in\bar{Z}}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)$ (13)
$\underset{\phantom{n\to\infty}}{=}\frac{|Z|}{n(n-1)}\left(\frac{1}{|Z|}\sum_{(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in Z}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)+\frac{|\bar{Z}|}{n(n-1)}\left(\frac{1}{|\bar{Z}|}\sum_{(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in\bar{Z}}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)$ (14)
=$\underset{n\to\infty}{=}\frac{1}{k}\cdot\mathbb{E}_{p(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-}\mid(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in Z)}[S]+\left(1-\frac{1}{k}\right)\cdot\mathbb{E}_{p(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-}\mid(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in\bar{Z})}[S]$ (15)
@@
We performed a recovery check analysis using a single layer from the iCNN method in Fig 11. The iCNN method generates an image by optimizing pixel values to make the image’s latent features similar to the target latent features [29]. In the pixel optimization condition (the left columns of each recovery image in Fig 11), we directly optimized the pixel values of images to minimize the mean squared loss between the latent features of the image as well as the total-variance (TV) loss of pixel values [65].
Additionally, the iCNN method can incorporate image generator networks (middle and right columns of each recovery image in Fig 11) to add constraints on image statistics. Instead of optimizing the pixel values, we optimized the parameters related to the generator networks to minimize the mean squared loss between the latent features obtained through the generator networks and the target latent features. As a weak image prior, we used Deep Image Prior (DIP) [68]. DIP utilizes a hierarchical U-Net architecture as an inherent prior for image tasks, capturing the statistical regularities of images without relying on a specific dataset. This model works effectively by optimizing a randomly initialized neural network that can be used as an image prior in various inverse problems such as denoising, super-resolution, and inpainting tasks. In our analysis, DIP started with a U-Net initialized with random noise. Subsequently, the latent features and weight parameters of DIP were optimized to minimize the difference between the network’s output and target DNN features. For the pre-trained image prior [69], we used the same generator model as in Shen et al. [29] optimizing the latent features of the pre-trained networks.

### v5

@@ 3 Discussion
Additionally, the independence of test stimuli from the model training processes requires careful examination, particularly when using pre-trained deep neural network (DNN) models and foundation models like CLIP or diffusion models. These models are typically trained on vast amounts of data available on the internet (Radford et al., 2021; Brown et al., 2020; Rombach et al., 2022), which likely includes public datasets such as MS-COCO (Lin et al., 2014) that are used in the NSD. This overlap raises potential concerns about the true independence of test stimuli, as we cannot rule out the possibility that these pre-trained models have acquired representations specifically tailored to the stimuli used for the model training. To address these issues and ensure a more rigorous evaluation of model generalization capabilities, researchers may consider using test stimuli that are not publicly available on the internet. This could include self-created stimuli or carefully curated datasets that have not been used in the training of widely used AI models (Shen et al., 2019b; Cheng et al., 2023). Such an approach would provide a more stringent test of a model’s ability to generalize to truly novel inputs.
Although we highlighted how stimulus overlap and limited diversity can lead to an overestimation of visual image reconstruction performance, such considerations do not apply to all reconstruction tasks. The dimensionality of the target space varies depending on the domain we seek to reconstruct; for instance, movement reconstruction often involves low-dimensional outputs (e.g., movement direction or velocity), where a small number of brain activity samples may suffice. In such cases, robust reconstruction is achievable by memorizing the brain-target pairs that cover the output space and classifying the test brain data into trained targets or interpolating between them. However, in visual image reconstruction with high-dimensional output space, evaluating zero-shot prediction is essential. Models must demonstrate true generalization to unseen stimuli rather than reflecting training data biases for establishing both reliability and practical utility.
@@
One of the remaining challenges in visual image reconstruction is the development of metrics for evaluating the quality and accuracy of the reconstructed images. The first and most critical step in assessing reconstruction results is to confirm a qualitative similarity between the reconstructed images and the perceived images through visual inspection across a diverse range of test sets. Following this, quantitative metrics should be employed for a more objective, high-throughput evaluation. However, as our analysis has suggested, it can be misleading to evaluate reconstruction by heavily relying on identification performance based on the relative similarity among alternatives (Koide-Majima et al., 2024). Even in cases where the reconstructed images only capture superficial information, such as categories or overall brightness, identification metrics can still be high. While identification performance can provide a useful benchmark, it should not be the sole metric for evaluating reconstruction quality. It is crucial to develop more appropriate similarity metrics that can accurately measure the perceptual similarity between the reconstructed and original images. One promising approach is to leverage image quality assessment (IQA) techniques from the computer vision field (Fu et al., 2024; Ding et al., 2020). These techniques are designed to quantify the perceptual quality of images and can be adapted to the specific requirements of visual image reconstruction.
@@
Our findings have profound implications for research integrity and responsible dissemination of scientific results. Visual image reconstruction methods have gained attention not only from neuroscientists but also from the general public and policymakers, sparking discussions about their potential applications and risks (UNESCO, 2023). These stakeholders often contemplate the possibilities of seamless information communication through the brain, such as in brain-machine interfaces (BMIs), or the dangers of unauthorized access to private information from brain activity. This interest may stem from the perception that brain activity data can be obtained easily and reliably in real-time. However, current technology and analysis methods fall short of these expectations. Beyond the limitations discussed in this study, there are additional challenges in the field. Most reconstruction methods analyze previously acquired brain data offline. The brain data used for reconstructing images are often averaged over multiple presentations of the test image, with only a few studies demonstrating single-trial reconstruction results (Miyawaki et al., 2008; Cheng et al., 2023). Further, it has been argued that subject cooperation is essential for reliably training and testing decoding models (Tang et al., 2023). Given these realities, public expectations often exceed current capabilities, and meeting these high demands in the short term is challenging. By clearly articulating these constraints, we can help manage expectations, prevent disappointment, and guide governments and companies away from misguided decisions. It is crucial to resist making overly optimistic claims about the ability to reconstruct arbitrary images. In light of these challenges, while the field of visual image reconstruction from brain activity holds great promise, it is our responsibility as researchers to ensure that its current capabilities and limitations are accurately communicated to all stakeholders.
3.1 Recommendations
Finally, we present several guidelines for critically testing visual image reconstruction methods. These suggestions build upon the limitations and challenges identified in our study and provide a pathway for future improvements in reconstruction research.
3.1.1 Stimulus design and data splits
Expand and control diversity. Collect or curate training datasets that span sufficient axes in the feature space so that new (unseen) visual images can be predicted. When possible, include artificial or carefully-designed stimuli as well as natural stimuli in the test set to provide clearer interpretability and control. Natural image reconstruction is not necessarily the ultimate goal, given that humans perceive both natural and artificial images.
Avoid overlaps in training and test stimuli. To evaluate true zero-shot capacity, ensure that test images do not overlap semantically or visually with training images. Identifying and removing near-duplicates or highly similar images from the test set helps prevent hidden “shortcut” solutions in which the model simply memorizes or classifies into known stimuli.
Use multiple and independent test sets. Consider separate test sets with varying complexity—e.g., natural images, artificial shapes, and out-of-distribution samples—to comprehensively assess generalizability. Disentangling performance across these varied sets can reveal whether a method is genuinely reconstructing novel content or only handling a narrow range of stimuli.
3.1.2 Model specification and latent feature choice
Confirm the generator’s recovery capability. Perform “recovery checks” by feeding true latent features (extracted directly from the original images) into the image generator. If the generator fails to reproduce the original images faithfully, it cannot serve as a valid reconstruction module. This step clarifies whether errors in the test phase stem from the latent translator or from a generator prone to hallucinations.
Use compositional, image-preserving features. Favor latent features (such as mid-level DNN layers) that retain sufficient image-level detail and compositional representations. This ensures that, with perfect translation from the brain, the original image can be reconstructed accurately. In contrast, purely semantic or text-based features often discard important visual details, limiting reconstruction fidelity.
Mitigate output dimension collapse. Choose or design translators (e.g., modular or sparse voxel selection approaches) to reduce collapse onto limited training-set clusters. Avoid overfitting to narrow categories by ensuring that the model maintains high-dimensional predictive capacity across all important feature dimensions.
3.1.3 Evaluation metrics and result transparency
Prioritize perceptual resemblance across diverse targets. Before focusing on quantitative metrics, visually confirm that reconstructions capture the perceptual features of all target stimuli. Testing fidelity on diverse or out-of-distribution samples is crucial for confirming genuine reconstruction rather than mere classification or retrieval. Avoid overemphasizing photorealism, as it can mask inaccuracies. Include extensive examples so readers can visually assess quality and variation.
Avoid cherry-picking and post-hoc selection. Generative models can produce multiple plausible outputs from a single latent feature. Selecting only the best-looking or most accurate images artificially inflates performance estimates. Present results transparently (e.g., showing random draws or evaluating robustness across different seeds) to offer a fair depiction of each model’s true reliability.
Use robust metrics beyond pairwise identification. Pairwise identification is easy to implement, but can overestimate performance especially in categorically structured data. Carefully design the selection of candidates when conducting identification analysis. In addition, supplement it with more stringent evaluation, such as fine-grained semantic checks, and advanced image-quality metrics (e.g., SSIM, DreamSim, other learned perceptual measures). Distinguish clearly between high-level semantic alignment and perceptual similarity.
3.1.4 Collaboration and ethical communication
Interdisciplinary collaborations. Close collaboration among neuroscientists, machine learning researchers, and cognitive scientists is vital for designing robust experiments, interpreting results correctly, and addressing complex technical pitfalls (e.g., data leakage, improper splits, or hallucinations by diffusion models).
Transparent reporting and data-sharing. Provide open-source code, clearly document training–test splits, and release relevant stimuli annotations to enable reproducibility. Transparency fosters collective progress, allowing others to replicate or extend your findings under more controlled or diverse conditions.
Realistic public and policy discourse. Communicate clearly that current reconstruction methods do not equate to unconstrained “mind reading” and often depend on carefully curated data. Highlight the role of subject cooperation, offline averaging, and limited generalizability so that stakeholders—such as policymakers, journalists, and the public—avoid overestimating immediate real-world capabilities.
In sum, authentic visual image reconstruction from brain activity requires careful management of dataset diversity and overlap, prudent model specification (especially in latent feature selection), and rigorous evaluation metrics beyond simple identification. By adhering to these recommendations, researchers can reduce the risk of reporting spurious reconstructions, bringing us closer to methods that genuinely reflect an individual’s perceptual experience.
4 Conclusions
Our critical analysis of recent generative AI-based visual image reconstruction methods revealed several limitations and challenges. We demonstrated that the apparent success of text-guided reconstruction methods primarily stems from a combination of classification into trained categories and hallucination through text-to-image diffusion, rather than genuine zero-shot reconstruction of novel images, which was the original goal of reconstruction studies. Our formal analysis revealed that predicting features with limited diversity can lead to output dimension collapse, where predictions become confined to patterns similar to the training set. Our simulation analysis demonstrated that successful zero-shot prediction requires training data with sufficient diversity to span the effective dimensions of the target feature space. We also pointed out that standard identification metrics can be misleading, especially when the target set has an underlying similarity structure. Additionally, we provided evidence that much of the input information is preserved at almost all hierarchical levels of deep neural networks. Finally, we pointed out that recent realistic reconstructions produced by generative AI models may appear convincing but do not necessarily reflect accurate representations of perceived visual experiences. These findings emphasize the need for more rigorous evaluation methods, diverse datasets, and careful interpretation of results in visual image reconstruction research. Future work should focus on prioritizing accurate reconstruction rather than naturalistic appearance. This objective would be achieved by utilizing compositional representations that can effectively span the feature space while maintaining the ability to recover original stimuli from latent features. As the field continues to attract attention from both researchers and the public, our results have important implications for research integrity and responsible development of neurotechnology, highlighting the need to balance scientific advancement with realistic expectations about current technological capabilities.
Methods
@@
We utilized two datasets: the Natural Scene Dataset (NSD; Allen et al., 2022) and the Deeprecon dataset (Shen et al., 2019b). Both datasets comprise visual stimuli and corresponding fMRI activity collected when subjects perceived the stimuli. In the NSD dataset, eight subjects were presented with MSCOCO images (Lin et al., 2014), yielding $30,000$ brain activity samples per subject, which is three times the amount provided by the Deeprecon dataset. The Deeprecon dataset includes fMRI activity data from subjects presented with both ImageNet images (Deng et al., 2009) and artificial images. It contains roughly $8,000$ brain samples per subject. Since this dataset is designed to evaluate reconstruction performance, the test stimuli were carefully selected. The test natural images were selected from ImageNet, which were in categories different from those used in the training. The artificial images were only used as test data to check the generalizability performance of the proposed reconstruction methods. In both datasets, we adopted the training–test split used in previous studies and utilized data from the first subject (S1 in the NSD and Subject 1 in the Deeprecon). Text-guided reconstruction methods require text annotations of images. For the NSD, text annotations accompanying the MSCOCO database were used. For the Deeprecon dataset, we collected text annotations for each experimental stimulus via crowd workers on Amazon Mechanical Turk, yielding five annotations per image. The text annotations of training stimuli are publicly available in the GitHub repository (https://github.com/KamitaniLab/GOD_stimuli_annotations).
@@
We utilized three image reconstruction methods: StableDiffusionReconstruction (Takagi and Nishimoto, 2023a), Brain-Diffuser (Ozcelik and VanRullen, 2023), MindEye2 (Scotti et al., 2024), and iCNN (Shen et al., 2019b). Each method employs two common steps: first, translating brain activity patterns into latent features of the stimuli, and second, generating images from these latent features using an image generator (Fig. 1). In the StableDiffusionReconstruction method, the latent features are the VAE (Kingma and Welling, 2014) features calculated from stimulus images and the CLIP text features (Radford et al., 2021) from the image annotations. The generator is StableDiffusion (Rombach et al., 2022). They first generate low-resolution images from the translated VAE features, and those images are further fed into the StableDiffusion model with translated text features to generate images. The generated images are regarded as reconstructed images from brain activity. In the Brain-Diffuser method, the latent features are the VDVAE (Child, 2021), CLIP vision features from stimulus images and CLIP text features from the text annotations of the image. The generator is Versatile diffusion (Xu et al., 2022). Similar to StableDiffusionReconstruction, low-resolution images are first generated from the translated VDVAE features, and these images are further used for the input of the versatile diffusion model with the translated vision and text features. The generated images are regarded as reconstructed images from brain activity. In the MindEye2 methods, the latent features are the variant of CLIP vision model features (OpenCLIP ViT/bigG-14). The generator is multiple Stable diffusion XL (SDXL) models (Podell et al., 2023). They first generate images from translated OpenCLIP ViT/bigG-14 features by unCLIP technique (Ramesh et al., 2022; Scotti et al., 2024). They then generate the final reconstruction by SDXL, integrating the generated images and text caption predicted from the translated OpenCLIP ViT/bigG-14 features by GiT Image2Text modules (Wang et al., 2022). In the iCNN method, the latent features are the intermediate output of the VGG19 layer (Simonyan and Zisserman, 2015). As a generator, they used the pre-trained image generator (Dosovitskiy and Brox, 2016), and they solved the optimization problem to minimize the discrepancy between the VGG19 features calculated from the generated images and the translated VGG19 features. Well-optimized images are regarded as reconstructed images.
@@
To investigate dataset diversity, we employed uniform manifold approximation and projection (UMAP), a nonlinear dimensionality reduction technique (McInnes et al., 2018), to learn a projection from a latent features space to a lower dimension (UMAP embedding space). We used both the training and test CLIP text features to learn the UMAP projection. These features were combined and standardized beforehand. The hyperparameters followed the official guide for clustering usage (https://umap-learn.readthedocs.io/en/latest/clustering.html) with cosine distance as a distance metric. The learned UMAP was also used to project the features predicted from brain activity (Fig. 6B). After standardizing the predicted features using the same mean and standard deviation parameters used in UMAP projection learning, we projected the predicted features onto the UMAP embedding space.
@@
We conducted a simulation analysis to illustrate output dimension collapse (Fig. 8) and to examine the generalization performance beyond the training data (Fig. 9). These analyses involve a teacher-student learning task. The teacher model generated the pairs of target features and brain activity as training samples, and the student model learned a mapping from the training samples. To imitate the feature translation situation from brain activity, observation noise was added to the brain data.
@@
For $N$ training samples, $X_{\mathrm{tr}}=[\mathbf{x}^{(1)},\ldots,\mathbf{x}^{(N)}]^{\top}$ and $Y_{\mathrm{tr}}=[\mathbf{y}^{(1)},\ldots,\mathbf{y}^{(N)}]^{\top}$, we trained the student model. by the ridge regression algorithm. The trained weight of ridge regression model $W$ can be calculated analytically:
@@
After training the student weight $W$, we illustrated the phenomenon of output dimension collapse as in Fig. 8 by predicting randomly generated data. First, we generated random test latent features from the Gaussian distribution with a mean of $\mathbf{0}$ and a variance equal to the training set variance scaled by a single scalar value ($9.0$). The corresponding test brain samples were obtained by translating the latent features by the teacher’s weight $\bar{A}$ and adding observation noise $\boldsymbol{\xi}$. The predicted latent features were derived by projecting the test brain samples by the learned student weight $W$. We set $D=512$, $\sigma_{\mathrm{intra}}^{2}=10/512$, $\sigma_{\mathrm{inter}}^{2}=100/512$, and $\sigma_{\mathrm{noise}}^{2}=10$.
We evaluated the zero-shot performance of the student model as in Fig. 9. We prepared two types of test samples: in-distribution test samples and out-of-distribution test samples. In-distribution test samples are generated from one of the (Gaussian) clusters used in training data:
@@
We mainly explored the dependency of the typical cluster identification performance on the following hyperparameters: dimension $D$, the number of clusters in the training set $C$, and the ratio of variance about the cluster structure $\sigma_{\mathrm{intra}}^{2}/\sigma_{\mathrm{inter}}^{2}$. We used a sufficiently large number of training samples $N=500,000$, and we kept $N$ constant while changing the above hyperparameters, especially $C$. Other parameters were also fixed in this simulation as follows: $\sigma_{\mathrm{intra}}^{2}+\sigma_{\mathrm{inter}}^{2}=110/D$, ${\bar{A}_{ij}}\sim\mathcal{N}(0,D^{-1/2})$, $\sigma_{\mathrm{noise}}^{2}=0.25$, $\lambda=1.0$, $n=100$ and $t=32$. We parameterized the scale of the variances ($\sigma_{\mathrm{intra}}^{2}$ and $\sigma_{\mathrm{inter}}^{2}$) and the teacher weights $\bar{A}_{ij}$ using the dimension $D$ so that the order of scale of the output samples were invariant from the dimension. When the number of clusters in the training set $C$ was less than $t=32$, we set $t=C$ instead.
@@
$\mathrm{Acc}$ $\underset{\phantom{n\to\infty}}{=}\frac{1}{n(n-1)}\left(\sum_{(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in Z}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)$
$\hskip 99.58464pt+\frac{1}{n(n-1)}\left(\sum_{(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in\bar{Z}}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)$ (13)
$\underset{\phantom{n\to\infty}}{=}\frac{|Z|}{n(n-1)}\left(\frac{1}{|Z|}\sum_{(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in Z}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)$
$\hskip 99.58464pt+\frac{|\bar{Z}|}{n(n-1)}\left(\frac{1}{|\bar{Z}|}\sum_{(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in\bar{Z}}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)$ (14)
=$\underset{n\to\infty}{=}\frac{1}{k}\cdot\mathbb{E}_{p(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-}\mid(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in Z)}[S]$
$\hskip 99.58464pt+\left(1-\frac{1}{k}\right)\cdot\mathbb{E}_{p(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-}\mid(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in\bar{Z})}[S]$ (15)
@@
We performed a recovery check analysis using a single layer from the iCNN method in Fig. 11. The iCNN method generates an image by optimizing pixel values to make the image’s latent features similar to the target latent features (Shen et al., 2019b). In the pixel optimization condition (the left columns of each recovery image in Fig. 11), we directly optimized the pixel values of images to minimize the mean squared loss between the latent features of the image as well as the total-variance (TV) loss of pixel values (Mahendran and Vedaldi, 2015).
Additionally, the iCNN method can incorporate image generator networks (middle and right columns of each recovery image in Fig. 11) to add constraints on image statistics. Instead of optimizing the pixel values, we optimized the parameters related to the generator networks to minimize the mean squared loss between the latent features obtained through the generator networks and the target latent features. As a weak image prior, we used Deep Image Prior (DIP; Ulyanov et al., 2018). DIP utilizes a hierarchical U-Net architecture as an inherent prior for image tasks, capturing the statistical regularities of images without relying on a specific dataset. This model works effectively by optimizing a randomly initialized neural network that can be used as an image prior in various inverse problems such as denoising, super-resolution, and inpainting tasks. In our analysis, DIP started with a U-Net initialized with random noise. Subsequently, the latent features and weight parameters of DIP were optimized to minimize the difference between the network’s output and target DNN features. For the pre-trained image prior (Dosovitskiy and Brox, 2016), we used the same generator model as in Shen et al. (2019b) optimizing the latent features of the pre-trained networks.